Find the electric potential difference Ve required to stop an electron (called a “stopping potential”) moving with an initial speed of 2.85 x 107 m/s. Would a proton travelling at the same speed require a greater or lesser magnitude of electric potential difference? Explain.


1
Expert's answer
2020-10-21T09:54:25-0400

Since the energy "W=U\times q" , where "U" is the potential difference, and "q" is the charge, and the kinetic energy "E=\frac{m\times v^2}{2}" , then "U\times q=\frac{m\times v^2}{2}" , hence "U=\frac{m\times v^2}{2\times q}=\frac{9.1\times 10^{-31}\times (2.85\times10^7)^2}{2\times 1.6\times 10^{-19}}\approx2309.84" V.

A proton will require a larger potential difference because it has a larger mass than an electron.